OpenStudy (anonymous):

What is the slope and y intercept for the graph of y-4x=-2

OpenStudy (anonymous):

1/2x i think

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

well you could rewrite this into y=mx+b form which would give you y=4x-2. This tells you that the slope is 4 (the number in front of the x). To find the y intercept you just need to plug in 0 for x, so you get y=4(0)-2 so y=-2 when x=0. The y intercept is (0,-2)

Nnesha (nnesha):

 you have to move -4x on right side so add 4x both side